OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. P(6, 2), Q(8, -4) _x_y__

OpenStudy (welshfella):

start with the 'point-slope' formula y - y1 = m(x - x1) where m = slope and (x1,y1) is a point on the lone to calculate m use rise/run =( y2-y1) / (x1-x1)

OpenStudy (anonymous):

could you please explain further?

OpenStudy (welshfella):

to work out m :- (x1,y1) = (6,2) (x2,y2) = (8,-4) plug these values into the rise/run formula to get the slope

OpenStudy (welshfella):